MARK WHICH STATEMENTS BELOW ARE TRUE, USING THE FOLLOWING, Consider Vf(x, y, z) in terms of a new coordinate system, x= x(u,A. The magnitude of a flow across a surface (the flux) is the magnitude of the flow vector per unit integrated over the area.
Answer #1

Sol: Given equations cura y x-xConw) y = y(u,vw) 4xx(u, viw) => 19) = x(s) î . y(s)j + 213) = arc length, hul I du = PartialA by The e magnitucle of flow across a surface as the magnitude of the flow vector per unit integrated over the area, - This